OpenStudy (anonymous):

find the horizontal aysmtope g(x)=9x^2-2x-5/2x^2-7x+2

OpenStudy (anonymous):

find the limit as x approaches infinity

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

divide both the numerator and the denominator by the term with the biggest degree, which is x^2 so you get g(x)=(9-(2/x)-(5/x^2))/(2-(7/x)+(2/x^2)) the limit of all the other terms containing x are 0. and so the limit of the function is g(x)=(9x^2-2x-5)/(2x^2-7x+2)=9/2